OpenStudy (dr0zier99):

Which test point holds true for y − 2x ≤ 1? (0, 2) (-2, 4) (1, 4) (5, 0)

OpenStudy (shamim):

Put x=0, y=2 y-2x=?

OpenStudy (dr0zier99):

2

OpenStudy (dr0zier99):

so its not a ok thanks

OpenStudy (shamim):

2 is not less than or equal to 1

OpenStudy (shamim):

Now try b option

OpenStudy (shamim):

Now i wanna know the result frm u

OpenStudy (dr0zier99):

the answer is d 5,0 y - 2x <= 1 0 - 2(5) <= 1 -10 <= 1